# Moments of inertia of rare-earth nuclei and the nuclear time-odd mean fields within exact solutions of the adiabatic theory

### v2

@@ Abstract
To quantitatively clarify the dependence of results on density functionals, we used seven commonly used Skyrme density functionals (SIII [30], SkM* [25], SkXce [31], SkO’ [32], SLy4 [33], UNEDF0 [34], and UNEDF1 [26]) with volume pairing to scrutinize the ratios ATDHFB/IB of rotational and axial vibrational inertia in 166Er, see Fig. 4. The ratios of inertia’s rotational and vibrational moments are linearly correlated with the isoscalar effective masses $m^{*}$ that characterize given functionals and decrease with slopes of --0.582 and --0.323, respectively. The correlations are quite strong with coefficients of determination ($R^{2}$) of 0.9027 and 0.9352. This shows that for these two collective modes, the impact of the time-odd fields on nuclear inertia proceeds through the isoscalar current terms $\bm{j}^{2}$ of the functionals, $\rho\tau-\bm{j}^{2}$ [30], which are linked to the effective-mass terms $\rho\tau$ by the gauge or Galilean invariance constraints. Nevertheless, even at $m^{*}/m$=1, the ratios ATDHFB/IB are larger than one, which shows that other time-odd mean fields also play a non-negligible role.
Figure 4: Correlation between the effective mass and the ratio between the ATDHFB and the IB rotational (a) and vibrational (b) moment of inertia of 166Er, calculated with various Skyrme functionals.
In summary, we proposed a novel iterative method to solve the ATDHFB equations exactly and we applied it to investigate the moment of inertia in deformed nuclei. For two Skyrme functionals SkM* and UNEDF1 and one Gogny functional D1S, we determined the rotational moment of inertia of the even-even nuclei between gadolinium and osmium. We obtained a good overall agreement with experimental data, although improvements in functional parameterizations, especially in the pairing channel, are required. The efficiency of our method bids well for the possibility of including the rotational moments of inertia in future adjustments of nuclear functionals.
@@ Acknowledgements
This work was partially supported by the STFC Grant Nos. ST/P003885/1 and  ST/V001035/1 and by the Polish National Science Centre under Contract No. 2018/31/ B/ST2/02220. We acknowledge the CSC-IT Center for Science Ltd., Finland, for the allocation of computational resources. This project was partly undertaken on the Viking Cluster, which is a high-performance computing facility provided by the University of York. We are grateful for computational support from the University of York High-Performance Computing service, Viking, and the Research Computing team.

### v3

@@ Abstract
To quantitatively clarify the dependence of results on density functionals, we used the seven commonly used Skyrme density functionals (SIII [30], SkM* [25], SkXce [31], SkO’ [32], SLy4 [33], UNEDF0 [34], and UNEDF1 [26]) with volume pairing to scrutinize the ratios ATDHFB/IB of rotational and axial vibrational collective inertia in 166Er, see Fig. 4. The rotational and vibrational collective inertia ratios are linearly correlated with the isoscalar effective masses $m^{*}$ that characterize given functionals and decrease with slopes of --0.582 and --0.323, respectively. The correlations are quite strong with coefficients of determination ($R^{2}$) of 0.9027 and 0.9352. This shows that for these two collective modes, the impact of the time-odd fields on nuclear collective inertia proceeds through the isoscalar current terms $\bm{j}^{2}$ of the functionals, $\rho\tau-\bm{j}^{2}$ [30], which are linked to the effective-mass terms $\rho\tau$ by the gauge or Galilean invariance constraints. Nevertheless, even at $m^{*}/m$=1, the ratios ATDHFB/IB are larger than one, which shows that other time-odd mean fields also play a non-negligible role.
Figure 4: Correlation between the effective mass and the ratio between the ATDHFB and the IB rotational (a) and vibrational (b) collective inertia in 166Er, calculated with various Skyrme functionals.
In summary, we proposed a novel iterative method to solve the ATDHFB equations exactly and we applied it to investigate the rotational moments of inertia in deformed nuclei. For two Skyrme functionals SkM* and UNEDF1 and one Gogny functional D1S, we determined the rotational moments of inertia of the even-even nuclei between gadolinium and osmium. We obtained a good overall agreement with experimental data, although improvements in functional parameterizations, especially in the pairing channel, are required. The efficiency of our method bids well for the possibility of including the rotational moments of inertia in future adjustments of nuclear functionals.
@@ Acknowledgements
This work was partially supported by the STFC Grant Nos. ST/P003885/1 and  ST/V001035/1 and by the Polish National Science Centre under Contract No. 2018/31/ B/ST2/02220. We acknowledge the CSC-IT Center for Science Ltd., Finland, for the allocation of computational resources. This project was partly undertaken on the Viking Cluster, which is a high-performance computing facility provided by the University of York. We are grateful for computational support from the University of York High-Performance Computing service, Viking, and the Research Computing team. We thank Grammarly® for its support with English writing.